OpenStudy (fredoneedshelp):

Translate into a multi-step equation Five more than the quotient of a number and 6 is 8. Use the variable x for the unknown number.

OpenStudy (ssddss02):

I have no clue soz m8

OpenStudy (fredoneedshelp):

oh

OpenStudy (fredoneedshelp):

@ellalaynemaguire

OpenStudy (ellalaynemaguire):

5+(x*6)=8

OpenStudy (ellalaynemaguire):

* is times

OpenStudy (fredoneedshelp):

ok

OpenStudy (fredoneedshelp):

it's sayin its incorrect

OpenStudy (ellalaynemaguire):

try 5+6x=8

OpenStudy (fredoneedshelp):

ok

OpenStudy (fredoneedshelp):

still wrong

OpenStudy (ellalaynemaguire):

5+(6x)=8

OpenStudy (ellalaynemaguire):

did that work?

OpenStudy (fredoneedshelp):

im checking

OpenStudy (fredoneedshelp):

nope it was x/5 +5 =8

OpenStudy (fredoneedshelp):

x/5 is a fraction

OpenStudy (fredoneedshelp):

Twice the difference of a number and 7 is equal to 8.

OpenStudy (fredoneedshelp):

use the variable y for the unknown number

OpenStudy (fredoneedshelp):

@MissSmartiez

OpenStudy (ellalaynemaguire):

2(y-7)=8 or 2(7-y)=8

OpenStudy (fredoneedshelp):

The sum of a number times 2 and 4 is equal to 8 . use c for the unknown number

OpenStudy (fredoneedshelp):

.